# Directional Invariants of Doubly Periodic Tangles

### v1

@@ Abstract.
In this paper we define novel topological invariants of doubly periodic tangles (DP tangles). DP tangles are embeddings of curves in the thickened plane that are repeated periodically in two directions. We first organize the components of a DP tangle into different interlinked compounds, which are invariants of a DP tangle. The notion of interlinked compound leads to the classification of DP tangles according to their directional type. We then prove that the directional type is an invariant of DP tangles using the concept of axis-motif, which can be viewed as the blueprint of a DP tangle.
@@
doubly periodic structures, tangles, thickened torus, motif, topological invariant, interlinked compound, directional element, directional type, axis-motif.
@@ 0. Introduction
Doubly periodic tangles (DP tangles) are complex entangled structures consisting of curves embedded in the thickened plane $\mathbb{E}^{2}\times I$ and can be viewed as lifts of links in the thickened torus, $T^{2}\times I$. DP tangles serve as a significant framework for analyzing and understanding the topological properties of interwoven filament systems across various scales, including but not limited to polymer networks [20, 21, 22], fabric-like structures [17, 7, 8, 16], and molecular chemistry [14, 15].
The topological classification of DP tangles is at least as hard a problem as the full classification of knots and links in the three-space. Such problems are approached by constructing topological invariants. We do that by first representing DP tangles by their generating cells, that we call motifs, which reduces the complexity of the classification problem. In this direction, several numerical, polynomial and finite type invariants have been constructed for doubly periodic textile structures (works initiated by Grishanov et al., cf. [9, 10, 11, 18, 12, 13, 1, 19]), which form a particular subclass of DP tangles. The equivalence relation that these invariants respect is based on assumptions of minimal motifs, that is motifs that are minimal for reproducing the DP tangles under 2-periodic boundary conditions. Note however that obtaining a minimal motif of a DP tangle is known to be a non trivial problem.
Motivated by the problem of the classification of DP tangles and the above works, we first established in [2] the topological equivalence of DP tangles on the level of arbitrary motifs of theirs, without the assumption of minimality (see Theorem 1.6).
@@ 1. Preliminaries on equivalence of DP tangles
In this section we recall the notions of DP tangles, their generating motifs and their equivalence, as established in [2]. The study in [2] is based on works initiated by Grishanov et al., cf. [9, 10, 11, 18]. It should be pointed out that these prior works focus on a subclass of DP tangles related to textiles (woven and knitted fabrics), and the results therein are based on assumptions of minimal motifs. In [2] we formulate and prove DP tangle equivalence for arbitrary DP tangles and arbitrary motifs of theirs.
@@ 1.1. DP tangles and motifs
Due to double periodicity of DP tangles, a DP tangle is defined as the lift of a link $\tau$ in the thickened torus $T^{2}\times I$ to the universal cover $\mathbb{E}^{2}\times I$, where $\mathbb{E}^{2}$ denote the Euclidean plane and $I=[0,1]$ the unit interval. An example is illustrated in Fig. 1. In particular, let $B=\{u,v\}$ be a basis of $\mathbb{E}^{2}$ and consider the covering map $\rho:\mathbb{E}^{2}\rightarrow{}T^{2}$ that assigns a longitude $l$ of $T^{2}$ to $u$ and a meridian $m$ of $T^{2}$ to $v$. Note that the covering map $\rho$ extends trivially to a covering map (also denoted) $\tilde{\rho}$: $\mathbb{E}^{2}\times I\rightarrow{}T^{2}\times I$.
@@
Figure 1. A link in $T^{2}\times I$, a corresponding flat motif and its DP tangle.
Definition 1.1 ([2]).
@@
Let $\tau$ be a link in $T^{2}\times I$. A doubly periodic tangle, or DP tangle, is the lift of $\tau$ under the covering map $\tilde{\rho}$, and is denoted by $\tau_{\infty}$. Moreover, a regular projection of $\tau$ onto $T^{2}\times\{0\}$ is called a (link) diagram of $\tau$, denoted by $d$, and the lift of $d$ under $\rho$ is called a doubly periodic diagram, or DP diagram, denoted by $d_{\infty}$. In this context, $d$ (or $\tau$) is called a motif for $d_{\infty}$ (resp. for $\tau_{\infty}$).
@@
Recall that a DP tangle comes equipped with a basis of $\mathbb{E}^{2}$ and a choice of a longitude-meridian pair $(l,m)$ for $T^{2}$. By a general position argument, $m$ and $l$ do not intersect crossings of $d$ and no arc of $d$ intersects a corner formed by $m$ and $l$. Moreover, the set of points $\Lambda(u,v)=\{xu+yv\,|\,x,y\in\mathbb{Z}\}$ generated by the basis $B$ of $\mathbb{E}^{2}$, defines a periodic lattice isotopic to $\mathbb{Z}^{2}$. Note that the flat torus $T^{2}$ containing a tangle diagram which represents the motif $d$ in $T^{2}$, that arises as the identification space, with respect to the boundary, of any one of the parallelograms of $\Lambda$, is called a flat motif. In particular, two bases $B=\{u,v\}$ and $B^{\prime}=\{u^{\prime},v^{\prime}\}$ generate the same point lattice $\Lambda=\Lambda(u,v)=\Lambda^{\prime}(u^{\prime},v^{\prime})$ if and only if for $x_{1},x_{2},x_{3},x_{4}\in\mathbb{Z}$,
$\begin{pmatrix}u^{\prime}\\ v^{\prime}\end{pmatrix}=\begin{pmatrix}x_{1}&x_{2}\\ x_{3}&x_{4}\end{pmatrix}\cdot\begin{pmatrix}u\\ v\end{pmatrix}\textit{ , where }\mid{x_{1}x_{4}-x_{2}x_{3}}\mid=\pm 1.$
@@
Figure 2. A DP diagram with a fixed lattice and two different bases of $\mathbb{E}^{2}$, leading to a shearing of the DP tangle.
@@
It is important to highlight that given a flat motif $d$ of a DP tangle $\tau_{\infty}$ associated to a lattice $\Lambda$, then any finite cover of $d$ also defines a flat motif of $\tau_{\infty}$ (see Fig. 3 for an example). We thus recall the definition of minimal lattice associated to a DP tangle.
Figure 3. Motif (a) is a stretching of (b); motif (d) is a contraction of (c); motifs (b) and (c) are scale equivalent.
Definition 1.2 ([2]).
@@
We define a minimal lattice of a DP tangle $\tau_{\infty}$, denoted by $\Lambda_{min}$, to be a lattice satisfying $\Lambda\subseteq\Lambda_{min}$, in the set of all periodic lattices $\Lambda$ of $d_{\infty}$, up to any area-preserving transformation of $\mathbb{E}^{2}$. Accordingly, the motif (resp. flat motif) $d=d_{\infty}/\Lambda_{min}$ is called a minimal (flat) motif for $d_{\infty}$.
@@ 1.2. Equivalence of DP tangles and their motifs
We first recall the definition of equivalence for DP tangles that preserve the double periodicity.
Definition 1.3 ([2]).
Two DP tangles (accordingly DP diagrams) are said to be equivalent if they can be obtained from each other by DP (diagrammatic) isotopies and orientation preserving invertible affine transformations of $\mathbb{E}^{2}$ carrying along the DP tangles (DP diagrams).
In [2], we studied DP tangle equivalence on the level of (flat) motifs along with the supporting lattices. DP isotopies are generated from local motif isotopies -namely surface isotopies and Reidemeister moves- and from torus homeomorphisms, applied on any motif of the DP tangle. Recall also that invertible affine transformations of the plane induce global isotopies of DP tangles, and they include re-scalings (stretches, contractions) -which are not area preserving transformations-, translations and rotations of the plane, as well as shear deformations (which are area preserving). DP tangle equivalence comprises also static changes, such as shifts or re-scalings of the underlying lattice (as, for example, motif duplication). For the purpose of this paper, we must recall two important notions of equivalence, namely Dehn equivalence and scale equivalence. First consider the lattice $\Lambda$, which can be generated by the two bases $B$ and $B^{\prime}$ defined above. Let us, now, assign the longitude $l$ of $T^{2}$ to $u^{\prime}$ and the meridian $m$ to $v^{\prime}$ of the basis $B^{\prime}$. This new covering map, say $\rho^{\prime}$, creates a new motif $d^{\prime}$, which differs from the motif $d$ associated to the basis $B$ by a finite sequence of Dehn twists, that are orientation preserving self-homeomorphisms of the torus, as explained in [9, 10] (for an illustration see Figure 4).
Figure 4. A longitudinal and a meridional Dehn twist of the torus.
Definition 1.4 ([2]).
Two (flat) motifs are said to be Dehn equivalent if they are related by a finite sequence of Dehn twists.
Next, considering now the equivalence between two distinct finite covers of the same motif, which thus lead to the same DP tangle, we have defined the following.
Definition 1.5 ([2]).
Let $d_{\infty}$ be a DP diagram and let $\Lambda_{0}$, $\Lambda_{1}$ and $\Lambda_{2}$ be three (not necessarily distinct) point lattices such that $\Lambda_{1}\subseteq\Lambda_{0}$ and $\Lambda_{2}\subseteq\Lambda_{0}$. Moreover, let $d_{0}=d_{\infty}/\Lambda_{0}$, $d_{1}=d_{\infty}/\Lambda_{1}$ and $d_{2}=d_{\infty}/\Lambda_{2}$ be flat motifs of $d_{\infty}$. Then, $d_{1}$ and $d_{2}$ arise as adjacent copies of $d_{0}$, according to the inclusion relations of the lattices. Then $d_{1}$ and $d_{2}$ are said to be scale equivalent. The notion of scale equivalence extends also to the corresponding motifs, which realize different finite covers of $T^{2}$.
These notions lead to the following generalized Reidemeister theorem for DP tangles, proved in [2], that is very important to define the notion of invariant of DP tangles.
@@
Figure 5. Different generating flat motifs for the same DP diagram and transformations among them, culminating to the minimal motif (d).
@@ 2.1. The class of an interlinked compound
Let $\tau$ be a (flat) motif in $T^{2}\times I$ and $\mu$ be a set of closed components of $\tau$ such that,
@@
Then, $\mu$ is said to be an interlinked compound of $\tau$. The definition carries through, analogously, to the DP tangle $\tau_{\infty}$.
Note that Definition 2.1 allows for an interlinked compound to be just a single component of the motif. In Figure 6, the motifs (a) and (e) consist of two interlinked components, the motifs (b) and (h) of three interlinked components, while the motifs (c), (d), (f) and (g) contain only one interlinked compound. Similarly, in Figure 1 we have one interlinked compound in the motif and the flat motif, which lifts to an infinitum of identical interlinked compounds in $\tau_{\infty}$.
Figure 6. Examples of flat motifs.
@@
Figure 7. Examples of flat motifs consisting of a single interlinked compound.
@@
Figure 7 (a) shows an example of a null-homotopic compound. Figure 7 (b) – (h) are examples of ribbon compounds, while Figure 7 (h) – (y) are examples of cover compounds.
@@ 2.2. Subclasses of interlinked compounds
From all possibilities of the structure of an interlinked compound we will single out and name some cases of interest. More precisely, let $\mu$ be an interlinked compound of a motif $\tau$ resp. of the DP tangle $\tau_{\infty}$.
@@ 3.1. The notion of directions
Let $\tau$ be a (flat) motif. An isolated knot in $\tau$ is a component of a null-homotopic compound. This compound may be interlinked with compounds of other classes in $\tau$.
@@ 3.2. Axis-motifs
    Motif (a) contains two isolated knots and three essential components of the same direction. So, its direction is the set $\{(0,0),(1,0)\}$ of 2 distinct elements, and its corresponding axis-motif (${}^{\prime}$) contains two trivial knots and a $(3,0)$-torus link.
    Motif (b) contains three disjoint ribbon compounds; the leftmost consists of two essential components, the middle one contains three essential component and the rightmost is a chain-link. The three compounds have the same direction, thus the motif has direction the singleton $\{(0,1)\}$, and its corresponding axis-motif (${}^{\prime}$) is a $(0,6)$-torus link.
    Motif (c) contains a null-essential ribbon compound made of two essential components and three isolated knots. So, it has direction $\{(0,0),(2,1)\}$ and its corresponding axis-motif (${}^{\prime}$) contains a $(4,2)$-torus link and three trivial knots.
    Motif (d) contains an essential cover compound made of three essential components. It has direction $\{(1,0),(1,2),(-1,2)\}$ and its corresponding axis-motif (${}^{\prime}$) contains a $(1,0)$-torus knot, a $(1,2)$-torus knot and a $(-1,2)$-torus knot.
    Motif (e) contains a chain-polycatenane cover compound made of two interlinked chain-links, and a null-homotopic compound consisting of a single knot. It has direction $\{(0,0),(1,0),(0,1)\}$ and its corresponding axis-motif (${}^{\prime}$) contains a $(1,0)$-torus knot, a $(0,1)$-torus knot and a trivial knot.
@@
    Motif (h) contains a cover compound consisting of an interlinkage of: a full-polycatenane of one knot, a chain-link, three essential curves (pink, green, black), and a trivial knot. It also contains two null-homotopic compounds of one and two components respectively. It has direction $\{(\infty,\infty),(0,0),(1,0),(0,1)\}$ and its corresponding axis-motif (${}^{\prime}$) contains a $(2,0)$-torus link, a $(0,2)$-torus link, four trivial knots, and one non-contractible trivial knot.
Figure 8. The corresponding axis-motifs of the flat motifs in Figure 6.

### v2

@@ Abstract.
In this paper we define novel topological invariants of doubly periodic tangles (DP tangles). DP tangles are embeddings of curves in the thickened plane with translational symmetries in two independent directions. We first organize the components of a DP tangle into different interlinked compounds, which are invariants of a DP tangle. The notion of interlinked compound leads to the classification of DP tangles according to their directional type. We then prove that the directional type is an invariant of DP tangles using the concept of axis-motif, which can be viewed as the blueprint of a DP tangle.
@@
doubly periodic structures, translational symmetry, tangles, thickened torus, motif, topological invariant, interlinked compound, directional element, directional type, axis-motif.
@@ 0. Introduction
A doubly periodic tangle (DP tangle) is an embedding of curves in the thickened plane $\mathbb{E}^{2}\times I$ which is symmetric under translations in two transverse directions. So, DP tangles can be viewed as lifts of links in the thickened torus, $T^{2}\times I$. An example can be viewed in Figure 1. DP tangles serve as a significant framework for analyzing and understanding the topological properties of interwoven filament systems across micro-, meso- and macro-scales, including, but not limited to, polymer melts [21, 22, 23], fabric-like structures [18, 8], molecular chemistry [16, 17], and cosmic filaments [1, 14, 15].
The topological classification of DP tangles is at least as hard a problem as the full classification of knots and links in the three-space. Such problems are approached by constructing topological invariants. We do that by first representing a DP tangle by its quotient, that we call motif, under a periodic lattice, which serves as a frame for the double translational symmetry of the DP tangle. This approach reduces the complexity of the classification problem. In this direction, several numerical, polynomial and finite type invariants have been constructed for doubly periodic textile structures,such as woven and knitted fabrics (works initiated by Grishanov et al., cf. [9, 10, 11, 19, 12, 13, 2, 20]), which form a particular subclass of DP tangles. The equivalence relation that these invariants respect is based on assumptions of minimal motifs, that is motifs that are minimal for reproducing the DP tangles under 2-periodic boundary conditions. Note however that obtaining a minimal motif of a DP tangle is known to be a non-trivial problem [9].
Motivated by the classification problem of DP tangles and the above works, we first established in [3] the topological equivalence of DP tangles on the level of arbitrary motifs of theirs, without the assumption of minimality (see Theorem 1.4).
@@ 1. Preliminaries on equivalence of DP tangles
In this section we recall the notions of DP tangles, their generating motifs and their equivalence, as established in [3], where we formulate and prove DP tangle equivalence for arbitrary DP tangles and arbitrary motifs of theirs. This study is based on works initiated by Grishanov et al., cf. [9, 10, 11, 19]. It should be pointed out that these prior works focus on a subclass of DP tangles related to textiles (such as woven and knitted fabrics), and the results therein are based on assumptions of minimal motifs. In DP textiles, their motifs consist of only essential simple closed curves embedded in the thickened torus. One of the reasons is that the weaving and knitting machines that make these textiles cannot build null-homotopic components. In this study we consider the full theoretical generality.
@@ 1.1. DP tangles and motifs
Due to the double translational symmetry of DP tangles, a DP tangle is defined as the lift to $\mathbb{E}^{2}\times I$ of a link $\tau$ in the thickened torus $T^{2}\times I$, where $\mathbb{E}^{2}$ stands for the Euclidean plane and $I=[0,1]$ the unit interval. View Figure 1 for an illustrated example. More precisely, consider a basis $B=\{u,v\}$ of $\mathbb{E}^{2}$ and let $\rho:\mathbb{E}^{2}\rightarrow{}T^{2}$ be the covering map that assigns a meridian $m$ of $T^{2}$ to $v$ and a longitude $l$ of $T^{2}$ to $u$. This covering map $\rho$ extends to a covering map $\tilde{\rho}$: $\mathbb{E}^{2}\times I\rightarrow{}T^{2}\times I$.
@@
Figure 1. A DP tangle as the lift of a link in $T^{2}\times I$, and a corresponding flat motif.
Definition 1.1.
@@
A DP tangle, denoted $\tau_{\infty}$, is the lift under $\tilde{\rho}$ of a link $\tau$ in $T^{2}\times I$. Moreover, a DP diagram, denoted $d_{\infty}$, is the lift under $\rho$ of a diagram $d$ of $\tau$ in $T^{2}\times\{0\}$. We call $d$ (resp. $\tau$) a motif of $d_{\infty}$ (resp. $\tau_{\infty}$).
@@
Assigning further a basis $B=\{u,v\}$ of $\mathbb{E}^{2}$ to a DP tangle together with a choice of longitude $l$ and meridian $m$ for $T^{2}$, the set of points $\Lambda(u,v)=\{xu+yv\,|\,x,y\in\mathbb{Z}\}$ generated by $B$ defines a periodic lattice isomorphic to $\mathbb{Z}^{2}$. Moreover, the tangle diagram contained in the flat torus $T^{2}$ that arises as the identification space by identifying the opposite sides of the boundary of any one of the parallelograms of $\Lambda$, represents a motif $d$ in $T^{2}$ and is called a flat motif, also denoted $d$ and we shall write $d=d_{\infty}/\Lambda$. In particular, it is well-known that the same periodic lattice $\Lambda=\Lambda(u,v)=\Lambda^{\prime}(u^{\prime},v^{\prime})$ can be generated by two different bases $B=\{u,v\}$ and $B^{\prime}=\{u^{\prime},v^{\prime}\}$ if and only if for $x_{1},x_{2},x_{3},x_{4}\in\mathbb{Z}$,
[…]
@@
Figure 2. A fixed lattice for a DP tangle, with two different bases of $\mathbb{E}^{2}$ (green and blue), which give rise to a shearing of the DP tangle.
@@
It is important to highlight that given a flat motif $d$ of a DP tangle $\tau_{\infty}$ associated to a lattice $\Lambda$, then any finite cover of $d$ also defines a flat motif of $\tau_{\infty}$ (see Figure 3 for an example). We thus recall the definition of minimal lattice associated to a DP tangle.
Figure 3. Motif (a) is a stretching of (b); motifs (b) and (c) are scale equivalent; motif (d) is a contraction of (c).
Definition 1.2.
@@
A minimal lattice, denoted by $\Lambda_{min}$, of a DP tangle $\tau_{\infty}$, is defined as a lattice satisfying $\Lambda\subseteq\Lambda_{min}$, in the set of all periodic lattices $\Lambda$ of $d_{\infty}$, up to any area-preserving transformation of $\mathbb{E}^{2}$. Moreover, the (flat) motif $d_{min}=d_{\infty}/\Lambda_{min}$ is called a minimal (flat) motif of $d_{\infty}$.
@@ 1.2. Equivalence of DP tangles and their motifs
We first recall the definition of equivalence for DP tangles that preserves the double periodicity.
Definition 1.3.
Two DP tangles (resp. DP diagrams) are equivalent if they differ by doubly periodic (diagrammatic) isotopies and orientation preserving invertible affine transformations of the plane $\mathbb{E}^{2}$ that carry along the DP tangles (resp. DP diagrams).
In [3], we studied the equivalence of DP tangles as reflected in their flat motifs. More precisely, doubly periodic isotopies include local motif isotopies (that is, surface isotopies and Reidemeister moves), as well as global isotopies, induced by invertible affine transformations of the plane. These consist of both non-area preserving transformations such as re-scalings (stretches, contractions), as well as area preserving transformations like rigid translations and rotations of the plane, or shear deformations. Re-scalings correspond to torus isotopies, like inflation or contractions, while shear deformations correspond to Dehn twists of the underlying torus, which are orientation preserving self-homeomorphisms of the torus, as related in [9, 10] (see Figure 4 for an example). More precisely, consider first the lattice $\Lambda$, which is generated by two bases $B$ and $B^{\prime}$ as defined above. We now assign the longitude $l$ of $T^{2}$ to $u^{\prime}$ and the meridian $m$ to $v^{\prime}$ of $B^{\prime}$. The corresponding covering map, say $\rho^{\prime}:\mathbb{E}^{2}\rightarrow{}T^{2}$, generates a new motif $d^{\prime}$, which differs from the motif $d$ associated to the basis $B$ by a finite sequence of Dehn twists. Two (flat) motifs related by a finite sequence of Dehn twists are said to be Dehn equivalent. Finally, in the DP tangle isotopy one should also include static changes, such as shifts or re-scalings of the underlying lattice. Re-scalings of the underlying lattice correspond to distinct finite covers of the same motif and we shall be referring to them as scale equivalence. More precisely, let $\Lambda_{0}$, $\Lambda_{1}$ and $\Lambda_{2}$ be three (not necessarily distinct) periodic lattices associated to a DP diagram $d_{\infty}$ such that $\Lambda_{1}\subseteq\Lambda_{0}$ and $\Lambda_{2}\subseteq\Lambda_{0}$. Let also $d_{0}=d_{\infty}/\Lambda_{0}$, $d_{1}=d_{\infty}/\Lambda_{1}$ and $d_{2}=d_{\infty}/\Lambda_{2}$ be three (flat) motifs of $d_{\infty}$. Then, $d_{1}$ and $d_{2}$ arise as finite covers of $d_{0}$ and are said to be scale equivalent.
Figure 4. Two Dehn twist of the torus: a longitudinal (purple) and a meridional one (blue)
The above lead to the following theorem for the DP tangle equivalence, which is very important for defining topological invariants of DP tangles.
@@
Figure 5. Equivalence transformations among different generating motifs for the same DP tangle, culminating to the minimal motif (d). Same colours of frames indicate the same lattices, related by area-preserving transformations.
@@ 2.1. The class of an interlinked compound
Let $\tau$ be a motif in $T^{2}\times I$ and $\mu$ be a set of closed components of $\tau$ such that,
@@
Then, $\mu$ is said to be an interlinked compound of $\tau$. The definition carries through, analogously, to the flat motif of $\tau$ and to the DP tangle $\tau_{\infty}$.
Note that Definition 2.1 allows for an interlinked compound to be just a single component of the motif. In Figure 6, the motifs (a) and (e) consist of two interlinked components, the motifs (b) and (h) consist of three interlinked components, while the motifs (c), (d), (f) and (g) consist of only one interlinked compound. Similarly, in Figure 1 we have one interlinked compound in the motif and the flat motif, which lifts to an infinitum of identical interlinked compounds in $\tau_{\infty}$.
Figure 6. Examples of flat motifs. Different shades help identify the different interlinked compounds in the motifs.
@@
Figure 7. Examples of flat motifs consisting of a single interlinked compound. Motif (a) is a null-homotopic compound, motifs (b)–(g) are ribbon compounds, and motifs (h)–(y) are cover compounds.
@@
Figure 7 (a) shows an example of a null-homotopic compound. Figure 7 (b) – (g) are examples of ribbon compounds, while Figure 7 (h) – (y) are examples of cover compounds.
@@ 2.2. Subclasses of interlinked compounds
As basic interlinked compounds we consider the following: null-homotopic compound, essential ribbon compound, chain-link ribbon compound, full-polycatenane compound (see descriptions below). All other interlinked compounds can be created as interlinkages of the basic types. From all possibilities of the structure of an interlinked compound we will single out and name some cases of interest. In the lists below, the basic interlinked compounds are marked in bold face. The namings will be in accordance to the stronger class of compound involved in the interlinkage.
Let $\mu$ be an interlinked compound of a motif $\tau$ resp. of the DP tangle $\tau_{\infty}$. We distinguish the cases:
@@ 3.1. The notion of directions
Let $\tau$ be a (flat) motif. An isolated knot in $\tau$ is a component of a null-homotopic compound. This compound could be an interlinked compound by itself or may be interlinked with basic compounds of other stronger classes in $\tau$, according to Definition 2.4 (recall Subsection 2.2).
For example, in motif (n) of Figure 7, the upper small trefoil is an isolated knot interlinked with a ribbon and a cover compound.
@@ 3.2. Axis-motifs
    Motif (a) contains two isolated knots and three essential components of the same direction. So, its direction is the set $\{(0,0),(1,0)\}$ of 2 distinct elements, and its corresponding axis-motif (a) contains two trivial knots and a $(3,0)$-torus link.
    Motif (b) contains three disjoint ribbon compounds; the leftmost consists of two essential components, the middle one contains three essential component and the rightmost is a chain-link. The three compounds have the same direction, thus the motif has direction the singleton $\{(0,1)\}$, and its corresponding axis-motif (b) is a $(0,6)$-torus link.
    Motif (c) contains a null-essential ribbon compound made of two essential components and three isolated knots. So, it has direction $\{(0,0),(2,1)\}$ and its corresponding axis-motif (c) contains a $(4,2)$-torus link and three trivial knots.
    Motif (d) contains an essential cover compound made of three essential components. It has direction $\{(1,0),(1,2),(-1,2)\}$ and its corresponding axis-motif (d) contains a $(1,0)$-torus knot, a $(1,2)$-torus knot and a $(-1,2)$-torus knot.
    Motif (e) contains a chain-polycatenane cover compound made of two interlinked chain-links, and a null-homotopic compound consisting of a single knot. It has direction $\{(0,0),(1,0),(0,1)\}$ and its corresponding axis-motif (e) contains a $(1,0)$-torus knot, a $(0,1)$-torus knot and a trivial knot.
@@
    Motif (h) contains a cover compound consisting of an interlinkage of: a full-polycatenane of one knot, a chain-link, three essential curves (pink, green, black), and a trivial knot. It also contains two null-homotopic compounds of one and two components respectively. It has direction $\{(\infty,\infty),(0,0),(1,0),(0,1)\}$ and its corresponding axis-motif (h) contains a $(2,0)$-torus link, a $(0,2)$-torus link, four trivial knots, and one non-contractible trivial knot.
Figure 8. The corresponding axis-motifs of the flat motifs in Figure 6. Color coding helps identify elements in the motif and their corresponding axes in the axis motif.